Twin brothers, Andy and Brian, can mow their grandparents lawn together in 40 min. Brian could mow the lawn by himself in 18 min more than it would take Andy. How long would it take each person to mow the lawn?
Answer by richwmiller(17219)

40/a+40/b=1
b=a+18
a=72
b=90
40/72+ 40/90=1
5/9+4/9=1